An object with a mass of 781 kg is accelerated by a force of 406 N. What is the acceleration it feels?
1
Expert's answer
2019-12-17T09:54:59-0500
"a=\frac{F}{m}"

"a=\frac{406}{781}=0.520\frac{m}{s^2}"